OpenStudy (legit_sarahh):

write the inverse of x=2x-3

OpenStudy (solomonzelman):

you mean y=2x-3 ?

OpenStudy (solomonzelman):

1) switch the x and y. 2) solve for y.

OpenStudy (legit_sarahh):

yeah ok thanks

OpenStudy (solomonzelman):

So, y=2x-3 do step one, you get x=2y-3 can you rearrange this to make it y= ... /

OpenStudy (legit_sarahh):

1.5x=y is the answer right?

OpenStudy (solomonzelman):

you are a little off. Lets do this together \(\LARGE\color{black}{ x=2y-3 }\) add 3 to both sides you get \(\LARGE\color{black}{ x+3=2y }\) divide each term by 2. you get?

OpenStudy (legit_sarahh):

x+1.5= y

OpenStudy (solomonzelman):